OpenStudy (anonymous):

Can someone show me how to graph this please?? y= 1/3 x+3

OpenStudy (anonymous):

is it \(y=\frac{1}{3}x+3\) ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

your best bet is to plot a few well chosen points for example, if \(x=0\) then \(y=3\) so \((0,3)\) is on the graph also if \(x=3\) then \(y=\frac{1}{3}\times 3+3=1+3=4\) so \((3,4)\) is on the graph also \((6,5)\) and \((9,6)\) etc

OpenStudy (anonymous):

as \(x\) increases by 3, \(y\) increases by 1, which is another way of saying the slope is \(\frac{1}{3}\)

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):



OpenStudy (anonymous):

plot the points, connect the dots

OpenStudy (anonymous):

\((0,3), (6,4), (9,5)\) plot those, then draw a line that connect them